OpenStudy (emmaleelooney):

help? anyone? The function y = -16t2 + 486 models the height y in feet of a stone t seconds after it is dropped from the edge of a vertical cliff. How long will it take the stone to hit the ground? Round to the nearest hundredth of a second. A. 11.02 seconds B. 0.25 seconds C. 7.79 seconds D. 5.51 seconds

rishavraj (rishavraj):

u just need to plug y = 0 and solve equation for t

rishavraj (rishavraj):

@emmaleelooney there??

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

\[ -16t^2 + 486 =0\] or \[16t^2=486\]

OpenStudy (misty1212):

divide by 16, take the square root of the result

OpenStudy (giitter):

First thing to do is solve T first. \[y=−16t^2+486\] ~~~~~~~~~~~~~~ The first step is to flip the equation. \[−16t^2+486=y\] ~~~~~~~~~~~~~~~ The we would have to Add -486 to both sides. ~~~~~~~~~~~~~~~~~~~~ \[−16t^2+486+−486=y+−486\] \[−16t^2=y−486 \] ~~~~~~~~~~~~~~~~~~~~~~~~~~~ After we added both of the sides, now we are going to divide both sides by \[-16\]. \[−16t^2 −16 = y−486 −16 t^2= −1 16 y+ 243 8\] Last step! Take square root. \[ t=√ −1 16 y+ 243 8 or t=−√ −1 16 y+ 243 8\] ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ Answer is- \[t=√ −1 16 y+ 243 8 or t=−√ −1 16 y+ 243 8 \] ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ I hoped I helped you with your question! Do what you love, love what you do! TYSM FOR POSTING THIS QUESTION AND LEAVING IT OPENED!!!!!!!!!!!! ~Glitter P.S. Pm me if you have any questions/concerns about my my answerrrr!